# Robust assessment of asymmetric division in colon cancer cells

### v2

@@ Abstract
Asymmetric partition of fate determinants during cell division is a hallmark of cell differentiation. Recent work suggested that such a mechanism is hijacked by cancer cells to increase both their phenotypic heterogeneity and plasticity and in turn their fitness. To quantify fluctuations in the partitioning of cellular elements, imaging-based approaches are used, whose accuracy is limited by the difficulty of detecting cell divisions. Our work addresses this gap proposing a general method based on high-throughput flow cytometry measurements coupled with a theoretical framework. We applied our method to a panel of both normal and cancerous human colon cells, showing that different kinds of colon adenocarcinoma cells display very distinct extents of fluctuations in their cytoplasm partition, explained by an asymmetric division of their size. To test the accuracy of our population-level protocol, we directly measure the inherited fractions of cellular elements from extensive time-lapses of live-cell laser scanning microscopy, finding excellent agreement across the cell types. Ultimately, our flow cytometry-based method promises to be accurate and easily applicable to a wide range of biological systems where the quantification of partition fluctuations would help accounting for the observed phenotypic heterogeneity and plasticity.
@@ I Introduction
Increasing evidence is vouching for a more ubiquitous presence of fluctuations in the partitioning of cellular elements. In fact, asymmetric segregation has been observed in non differentiating cell populations as different as bacteria Mushnikov2019 , yeast Yang2015 , and tumor cells Buss2024 ; Chao2024 . In bacterial populations, differential segregation of cellular components has been linked to antibiotic resistance Lindner2008 . Similarly, Katajisto et al. Katajisto2015 , studying mitochondrial partition in stem-like mammalian epithelial cells, found that asymmetric segregation of older mitochondria enables cells to protect themselves from aging by preventing the accumulation of misfolded proteins. In addition, asymmetric centrosome partition has been observed in stem cells from human neuroblastoma and colorectal cancer, controlled by polarity factors and influenced by the segregation of subcellular vesicles during cell division Izumi2012 . To quantify the strength of fluctuations at division, the common route goes through fluorescence microscopy measurements of either fixed or live cells, where fluorescent dyes are used as a proxy for the cellular element of interest. The partition statistics are then estimated by looking at the fraction of fluorescence intensity inherited by the two daughter cells DeyGuha2011 ; DeyGuha2015 . This approach is, in principle, very informative and a lot of work has been done in recent years to write theoretical models of stochastic gene expression beentjes2020exact ; jia2023coupling ; dessalles2020models ; thomas2018sources that can account for the coupling of noise sources and can even be extended to time-lapse microscopy data of single cell’s growing dynamics jia2021frequency ; however, relying on microscopy data implies the need of the (often manual) identification of hundreds of division events and in some cases even complex experimental set-ups and techniques where cells (usually bacteria) have to be followed for a great number of generations jia2021frequency . These are highly time-consuming procedures that dampen the possibility of a wide characterization of the partition statistics of different cell types.
@@
Our work shows that (i) there is an exact analytical expression linking the inherited cellular component distribution dynamic to the specifics of the partition process. Based on this result, we (ii) proposed an experimental protocol to tag with fluorescent dyes cellular components and measure the dynamics of fluorescent distribution, whose analysis allows accurate estimates of the degree of asymmetry in the partition process. Applying our method to a panel of normal and cancerous human colon cells, we found that (iii) different lines of colon adenocarcinoma display very distinct extent of fluctuations in cytoplasm partition, reflecting in (iv) different degrees of asymmetric division of their size.
Figure 1: Modeling partitioning noise at cell division. a) Schematic representation of the growth and division process of a mother cell. The cell undergoes an initial phase of duplication, where its internal elements are multiplied, followed by a division phase, where these elements are partitioned between the two daughter cells. One daughter inherits a fraction $p$ of the mother elements, while the other receives the remaining fraction $q=1-p$. b) (blue) Idealized behavior of the components number of a cell element over time, considering three different noise terms: fluctuations in compound counts during growth, uncertainty in the timing of division, and noise in the compound’s partition fraction. (red) Same description but considering only production and degradation processes. c) Microscopy images of a single cell division event for an HCT116 cell, whose cytoplasm is stained with Celltrace Yellow. d) Time evolution of the distribution of the components number of a cellular element for a population of cells subjected only to partitioning noise. The distribution used for the simulation is a Gaussian distribution with mean $\mu=\frac{1}{2}$ and standard deviation $\sigma=0.07$. e) Examples of partition distributions $\Pi$, with increasing coefficient of variation. f) Mean ($\mu_{g}$) and variance ($\sigma_{g}$) of the components number as a function of the generations, g, for the proliferation of a population subject to different partition noise distributions. Different colors correspond to distributions with varying coefficients of variation, as represented in the top panels. The dashed line represents the theoretical behavior obtained from the model. Dots are colored according to the distributions shown in panel e). g) Behavior of the mean (left) and standard deviation (right) of the simulated dynamics compared to the expected theoretical behavior for a proliferating population assuming a sizer division strategy (see SI).
@@ II Results
II.1 Modeling cell population dynamic in presence of partitioning noise
@@
In this section, we propose a theoretical framework that is able to describe the evolution of the fluorescence intensity of a cell population stained with live fluorescent marker, by uncovering its dependence on the underlying partitioning process to which single cells are subjected. The rationale is that, if cells follow a common division rule, knowing the shape of the mother distribution, it should be able to determine the daughters’ one. By iterating this computation, one can trace the whole proliferation. The specificity of our approach lies in its specific design for live fluorescent markers. The selected component is tagged through a staining procedure with fluorophores that persist in the cell and its daughter cells, as long as the targeted protein is not degraded.
@@
Monitoring the time evolution of the abundance of specific cellular elements over multiple divisions in single-cell experiments, it is observed Soltani2016 that the counts of a component have a qualitatively similar dynamic to those shown in Figure  1(b). Indeed, variations in the components number are determined by: (i) the production and degradation processes along the duration of the cell growth phase; and (ii) non identical apportioning of the components between the two daughter cells. Here, we will assume that the number of components varies only due to partition events and look for a quantitative description of the evolution of the element distribution. By neglecting variability in division times and the intercurrent production and degradation of components, we are reducing the dynamics to a progressive dilution. The stability of our model with respect to these assumptions will be studied in the following section and more in detail in the Supplementary Information. Note that it is possible to model the complete growth and division process (see for instance refs.  expexp ; miotto2023determining ; Scalise2024 ; jia2023coupling ; thomas2018sources ), but it is more difficult to separate the different sources of noise. Here, we start by considering a population of cells characterized by a certain initial distribution, $P(m)_{0}$, of element m. Each cell $i$ of the population divides into two daughter cells that we can label unambiguously as $2i$ and $12i+$ 1, which inherit respectively $m_{2i}$ and $m_{2i+1}$ components. The process, exactly at division, must conserve the total number of elements, i.e. $m_{i}=m_{2i}+m_{2i+1}$ and we call partitioning fraction the quantity $f=\frac{m_{2i}}{m_{i}}$. To model noise in partitioning, we assume that at each division a random value for $f$ is extracted from a probability distribution function, $\Pi(f)$. Given these assumptions, we can write the probability distribution of the number of components of a daughter cell $m$ after $g$ divisions from the mother $M$ as:
[…]
@@
where $P_{k}(M)$ is the components distribution at generation $k$ and it is given by the sum of the distributions of all siblings cells of the $k$-th generation. After $g$ generations, the cell lineage generated from a single cell forms a lineage tree composed of $n_{g}=2^{g}$ cells. Without losing generality, the division probability can be expressed, as:
@@
Therefore, we can write the probability distribution, $P(m)$, of the daughter cells inheriting a fraction f, of the mother elements:
[…]
As is often the case, it is convenient to characterize the distributions in terms of their moments miotto2023determining . We identify with $m_{2i}$ ($m_{2i+1}$) the subpopulation of cells inheriting a fraction $f$ ($1-f$). In particular, the mean of the daughter cell component distribution can be expressed, for $2i$, as :
[…]
@@
Where $E[\cdot]$ stands for the average. Note that every computation can be symmetrically done for the siblings subpopulation $12i+$ 1. Thus, we have that after a single generation, the mean number of inherited elements of the daughter cells depends on the asymmetry of the $\Pi(f)$ and on the mean of the $P(m_{i})$.
@@
The total distribution is the sum of $2^{g}$ sub-populations which derives from the division process. For example, if we assume $\Pi(f)=\frac{1}{2}(\delta(f)+\delta(1-f))$, after one division we would have a population distributed around $f\mu$ and one at $(1-f)\mu$. After two division the populations would grow to four sub-populations centered in $f^{2}\mu$, $f(1-f)\mu$, $(1-f)f\mu$ and $(1-f)^{2}\mu$ and so on.
@@
where $\mu_{g}^{k}$ is the mean of k-th sub-population at generation g and $2^{g}$ is the total number of sub-populations for that generation. This equation [5] can be simplified into the form (see Supplementary Note for the full computation):
@@
where $\mu_{0}$ is the mean of the initial population. It is relevant to notice, that independently of the characteristic of $\Pi(f)$, the mean always halves itself at each generation. This result is simply understood considering that for each cell that inherits a fraction $f^{\prime}$ of the compound, its sibling inherits a fraction $1-f^{\prime}$. Hence, at every division, the compound counts is diluted by an average factor of 2 in the population, and partition distributions with different properties lead to the same behavior of $\mu_{g}$ (see Figure 1(f)). Therefore, no relevant insights on the $\Pi(f)$ can be obtained from the first moment. We thus moved to consider the second moment, i.e. the variance. Similarly to what is done for the mean, the variance of the inherited fraction of elements can be expressed as:
@@
and in turn, the variance of a generation, starting from the consideration of the variance of a mixture of distributions can be written as (see Supplementary materials for details):
@@
It is possible to obtain a recursive equation for $A_{g}$ (see Supplementary Notes), $A_{g}=A_{g-1}E[f^{2}]=A_{0}E[f^{2}]^{g}$ which leads to a compact form for equation [8]:
@@
Finally, with equation 10 we can link the variance of the components distribution for the entire cell population at generation g, to the properties of $\Pi(f)$. It is interesting to highlight that there is no dependence on the bias asymmetry of the process. The distribution $\Pi(f)$ is necessarily symmetric, as daughter cells are indistinguishable. Therefore, by bias we refer to a situation where the most probable outcome of the division process is asymmetric. To this order, the only relevant feature shaping the dynamics is the extent of the fluctuations, regardless of their origin. Indeed, as shown in Figure 1(e), the same $\Pi(f)$, but with a different variance shows a different behavior for $\sigma_{g}^{2}$.
Our analytical results show that the dynamical behavior of the population variance depends non-trivially on the second moment of the underlining partitioning distribution. The next section will be devoted to demonstrating how this condition can be exploited in experiments to measure the partition noise component and to evaluate the stability of our assumptions with respect to concurrence of the other noise sources (see also Figure 1g).
@@ II.2 Measuring fluctuations via flow cytometry
As shown by Peruzzi et al. peruzzi2021asymmetric for leukemia cells growing in suspension, it is possible to use flow cytometry to follow the evolution of properly marked cellular elements in time. Here, we generalized the protocol considering a panel of colon cell lines, i.e. Caco2, a human epithelial colorectal adenocarcinoma cell line; HCT116, a human colorectal carcinoma cell line; and CCD-18Co, a normal human colon fibroblast cell line. To measure the degree of asymmetric division, we marked cells’ cytoplasm via a fluorescent dye and followed its dilution through successive generations. As explained in detail in the Materials and Methods section, the experimental procedure we used is based on the following main steps: staining of the selected compound, sorting of the initial population, plating and acquisition of the distinct populations in the following days. The sorting allows us to obtain an initial population with a narrow peak at a specific fluorescence intensity, and it plays a dual role (a full explanation can be found in the Supplementary Information): it indirectly induces cell synchronization and enables a clearer reconstruction of the coexisting generations.
To achieve this, we used fluorescence-activated cell sorting (FACS) to isolate a specific cell population based on cell morphology and fluorescence intensity. The sorted population is divided into distinct wells, one per acquisition, in equal numbers and kept under identical growth conditions to mitigate inoculum-dependent variability among samples  EnricoBena2021 . Figure 2(a) shows an example of the time course dynamics obtained with HCT116 cells (see Methods for details). Each distribution represents a different acquisition of a whole well at increasing times from plating (0 to 84 hours from the bottom to the top). It is possible to neatly observe the succession of peaks associated with different generations that succeed one another and the average decrease of the fluorescence intensity. Note that different generations can coexist at the same time point. One can spot the flow of cells towards higher generations and the behavior of the peak heights, which grow over time, reach dominance and then decay. The shift on the y-axis allows to follow the evolutionary dynamics while being able to see the overlapping of the same generation at different time points. As in each measurement, the distribution of fluorescent intensity $I$, is a mixture of distributions corresponding to different generations:
@@
To obtain the necessary information on the properties of each generations we performed a fitting of the data via a Gaussian Mixture Model (GMM) combined with an Expectation Maximization algorithm (details are explained in Materials and Methods). Indeed, assuming that each generation is log-normally distributed with mean $\mu_{g}$ and variance $\sigma_{g}^{2}$ of the corresponding Gaussian distribution, equation 11 can be rewritten as :
@@
An example of the outcome of the GMM algorithm is shown in Figure 2(b) for two acquisitions (48h and 60 h from plating). Blue-shaded distributions represent experimental data, while the best solution of the Gaussian mixture is shown with a solid line. Each generation’s fit is identified with a different color. Upon an overnight, we can identify generations 0, 1, and 2 in both the acquisitions. One can see that their relative fractions changed: generation 3 became dominant overnight and generation 4 appeared.
@@
Figure 2(c) shows the behaviors of $\mu_{g}$ and $\sigma_{g}$ as functions of the generation number $g$ for three independent replicas of the time course experiments with CaCo2 cells. Dashed lines represent the best fit of the data against Equation 5 for the mean and Equation 10, which instead yields the variance $\sigma_{f}$ of the partition distribution probability, $\Pi(f)$. The outcomes of the fitting procedure on the time course dynamic for all the replicas of three studied cell lines are reported in Figure 2(d). To evaluate the degree of asymmetry we have defined the division asymmetry as the percentage of the coefficient of variation of the obtained partition distribution (CV: $\frac{\sigma}{\mu})$. In particular, Caco2 and CCD18Co cells show a higher level of asymmetry with respect to HCT116. As already stated in the description of the theoretical model, we consider only partitioning noise and neglect other potential sources: (i) production and degradation processes, and (ii) variability in cell cycle length. Since we use non-endogenous live fluorescent dyes, production processes are excluded by definition, while degradation results in a negligible time-dependent decrease in mean fluorescence intensity as used dyes are optimized to maintain stable fluorescence levels for the duration of the experiment. Indeed, by looking at the data in Figure 2 (c) the values of the measured $\mu_{g}$ seem to only fluctuate around the theoretical line, but with the same slope, indicating that degradation can be neglected without major consequences. The effects of correlations between the inherited fractions of cellular components and the fluctuations in the duration of the cell cycle are in principle more difficult to disentangle. In this respect, it is widely known that cells adjust their growth/division strategy to mitigate size fluctuations at division miotto2024size ; thomas2018analysis . According to the ’sizer’ model, they divide upon reaching a certain size so that, smaller cells, i.e. those that inherit a smaller fraction of their mother’s volume, will have longer division times and vice-versa. This results in fluctuations in the cell division times that may be coupled with partitioning noise.
@@
To assess the robustness of our model under such conditions, we performed simulations assuming that cells follow a sizer-like division strategy, while the marked component partition perfectly matches cell size partition ( see Supplementary Information for more details). This scenario maximizes the coupling among the two noise sources. As can be observed in Figure 1(g), the behaviors of the mean value $\mu_{g}$ and the variance $\sigma_{g}$ deviate from the the case in which division time is uncoupled with partition noise. However, the values obtained are still within the estimated experimental error (see Supplementary Figure 3). It is worth noticing that the presence/absence of the increase in the mean, $\mu_{g}$, for newer generations may highlight the presence of the coupling between the two noise sources.

### v3

@@ Abstract
Asymmetric partition of fate determinants during cell division is a hallmark of cell differentiation. Recent work suggested that such a mechanism is hijacked by cancer cells to increase both their phenotypic heterogeneity and plasticity and, in turn, their fitness. To quantify fluctuations in the partitioning of cellular elements, imaging-based approaches are used, whose accuracy is limited by the difficulty of detecting cell divisions. Our work addresses this gap, proposing a general method based on high-throughput flow cytometry measurements coupled with a theoretical framework. We applied our method to a panel of both normal and cancerous human colon cells, showing that different kinds of colon adenocarcinoma cells display very distinct extents of fluctuations in their cytoplasm partition, explained by an asymmetric division of their size. To test the accuracy of our population-level protocol, we directly measure the inherited fractions of cellular elements from extensive time-lapses of live-cell laser scanning microscopy, finding excellent agreement across the cell types. Ultimately, our flow cytometry-based method promises to be accurate and easily applicable to a wide range of biological systems where the quantification of partition fluctuations would help account for the observed phenotypic heterogeneity and plasticity.
@@ I Introduction
Increasing evidence is vouching for a more ubiquitous presence of fluctuations in the partitioning of cellular elements. In fact, asymmetric segregation has been observed in non-differentiating cell populations as different as bacteria Mushnikov2019 , yeast Yang2015 , and tumor cells Buss2024 ; Chao2024 . In bacterial populations, differential segregation of cellular components has been linked to antibiotic resistance Lindner2008 . Similarly, Katajisto et al. Katajisto2015 , studying mitochondrial partition in stem-like mammalian epithelial cells, found that asymmetric segregation of older mitochondria enables cells to protect themselves from aging by preventing the accumulation of misfolded proteins. In addition, asymmetric centrosome partition has been observed in stem cells from human neuroblastoma and colorectal cancer, controlled by polarity factors and influenced by the segregation of subcellular vesicles during cell division Izumi2012 . To quantify the strength of fluctuations at division, the common route goes through fluorescence microscopy measurements of either fixed or live cells, where fluorescent dyes are used as a proxy for the cellular element of interest. The partition statistics are then estimated by looking at the fraction of fluorescence intensity inherited by the two daughter cells DeyGuha2011 ; DeyGuha2015 . This approach is, in principle, very informative and a lot of work has been done in recent years to write theoretical models of stochastic gene expression beentjes2020exact ; jia2023coupling ; dessalles2020models ; thomas2018sources that can account for the coupling of noise sources and can even be extended to time-lapse microscopy data of single cell’s growing dynamics jia2021frequency ; however, relying on microscopy data implies the need of the (often manual) identification of hundreds of division events and in some cases even complex experimental set-ups and techniques where cells (usually bacteria) have to be followed for a great number of generations jia2021frequency . These are highly time-consuming procedures that dampen the possibility of a wide characterization of the partition statistics of different cell types.
@@
Our work shows that (i) there is an exact analytical expression linking the inherited cellular component distribution dynamic to the specifics of the partition process. Based on this result, we (ii) proposed an experimental protocol to tag with fluorescent dyes cellular components and measure the dynamics of fluorescent distribution, whose analysis allows accurate estimates of the degree of asymmetry in the partition process. Applying our method to a panel of normal and cancerous human colon cells, we found that (iii) different lines of colon adenocarcinoma display very distinct extents of fluctuations in cytoplasm partition, reflecting in (iv) different degrees of asymmetric division of their size.
Figure 1: Modeling partitioning noise at cell division. a) Schematic representation of the growth and division process of a mother cell. The cell undergoes an initial phase of duplication, where its internal elements are multiplied, followed by a division phase, where these elements are partitioned between the two daughter cells. One daughter inherits a fraction $p$ of the mother’s elements, while the other receives the remaining fraction q=1pq=1-p. b) (blue) Idealized behavior of the components number of a cell element over time, considering three different noise terms: fluctuations in compound counts during growth, uncertainty in the timing of division, and noise in the compound’s partition fraction. (red) Same description, but considering only production and degradation processes. c) Microscopy images of a single cell division event for an HCT116 cell, whose cytoplasm is stained with Celltrace Yellow. d) Time evolution of the distribution of the components’ number of a cellular element for a population of cells subjected only to partitioning noise. The distribution used for the simulation is a Gaussian distribution with mean $\mu=\frac{1}{2}$ and standard deviation $\sigma=0.07$. e) Examples of partition distributions $\Pi$, with increasing coefficient of variation. f) Mean ($\mu_{g}$) and variance ($\sigma_{g}$) of the number of components as a function of the generations, g, for the proliferation of a population subject to different partition noise distributions. Different colors correspond to distributions with varying coefficients of variation, as represented in the top panels. The dashed line represents the theoretical behavior obtained from the model. Dots are colored according to the distributions shown in panel e). g) Behavior of the mean (left) and standard deviation (right) of the simulated dynamics compared to the expected theoretical behavior for a proliferating population assuming a sizer division strategy (see SI).
@@ II Results
II.1 Modeling cell population dynamics in the presence of partitioning noise
@@
In this section, we propose a theoretical framework that is able to describe the evolution of the fluorescence intensity of a cell population stained with a live fluorescent marker, by uncovering its dependence on the underlying partitioning process to which single cells are subjected. The rationale is that, if cells follow a common division rule, knowing the shape of the mother distribution, it should be able to determine the daughters’ one. By iterating this computation, one can trace the whole proliferation. The specificity of our approach lies in its specific design for live fluorescent markers. The selected component is tagged through a staining procedure with fluorophores that persist in the cell and its daughter cells, as long as the targeted protein is not degraded.
@@
Monitoring the time evolution of the abundance of specific cellular elements over multiple divisions in single-cell experiments, it is observed Soltani2016 that the counts of a component have a qualitatively similar dynamic to those shown in Figure 1(b). Indeed, variations in the number of components are determined by: (i) the production and degradation processes along the duration of the cell growth phase; and (ii) non-identical apportioning of the components between the two daughter cells. Here, we will assume that the number of components varies only due to partition events and look for a quantitative description of the evolution of the element distribution. By neglecting variability in division times and the intercurrent production and degradation of components, we are reducing the dynamics to a progressive dilution. The stability of our model with respect to these assumptions will be studied in the following section and more in detail in the Supplementary Information. Note that it is possible to model the complete growth and division process (see, for instance, refs.  expexp ; miotto2023determining ; Scalise2024 ; jia2023coupling ; thomas2018sources ), but it is more difficult to separate the different sources of noise. Here, we start by considering a population of cells characterized by a certain initial distribution, $P(m)_{0}$, of element m. Each cell $i$ of the population divides into two daughter cells that we can label unambiguously as $2i$ and $2i+1$, which inherit respectively $m_{2i}$ and $m_{2i+1}$ components. The process, exactly at division, must conserve the total number of elements, i.e. $m_{i}=m_{2i}+m_{2i+1}$, and we call the partitioning fraction the quantity $f=\frac{m_{2i}}{m_{i}}$. To model noise in partitioning, we assume that at each division a random value for $f$ is extracted from a probability distribution function, $\Pi(f)$. Given these assumptions, we can write the probability distribution of the number of components of a daughter cell $m$ after $g$ divisions from the mother $M$ as:
$P_{g}(m)=\left(\frac{1}{2^{g-1}}\right)\sum_{k=0}^{2^{g-1}}\int dMP(m|M)P^{k}_{g-1}(M)$ (1)
@@
where $P^{k}_{g-1}(M)$ are the components subdistributions present at generation g1g-1. After $g$ generations, the cell lineage generated from a single cell forms a lineage tree composed of $n_{g}=2^{g}$ cells. Without losing generality, the division probability can be expressed as:
@@
Therefore, we can write the probability distribution, $P(m)$, of the daughter cells inheriting a fraction $f$ of the mother elements:
$P(m)=\int dfdM\delta(m-fM)\Pi(f)P(M)$ (3)
As is often the case, it is convenient to characterize the distributions in terms of their moments miotto2023determining . We identify with $m_{2i}$ ($m_{2i+1}$) the subpopulation of cells inheriting a fraction $f$ (1f1-f). In particular, the mean of the daughter cell component distribution can be expressed, for $2i$, as :
$\mu_{2i}=E[m_{2i}]=\int P(m_{2i})m_{2i}\,dm_{2i}$
$=\int df\,dm_{i}\,dm_{2i}\,m_{2i}\delta(m_{2i}-fm_{i})\Pi(f)P(m_{i})$
@@
Where $E[\cdot]$ stands for the average. Note that every computation can be symmetrically done for the siblings subpopulation $2i+1$. Thus, we have that after a single generation, the mean number of inherited elements of the daughter cells depends on the asymmetry of the $\Pi(f)$ and on the mean of the $P(m_{i})$.
@@
The total distribution is the sum of $2^{g}$ sub-populations, which is derived from the division process. For example, if we assume $\Pi(f)=\frac{1}{2}(\delta(f)+\delta(1-f))$, after one division we would have a population distributed around $f\mu$ and one at $(1-f)\mu$. After two divisions, the populations would grow to four sub-populations centered in $f^{2}\mu$, $f(1-f)\mu$, $(1-f)f\mu$, and $(1-f)^{2}\mu$, and so on.
@@
where $\mu_{g}^{k}$ is the mean of the k-th sub-population at generation g and $2^{g}$ is the total number of sub-populations for that generation. Equation 5 can be simplified into the form (see Supplementary Note for the full computation):
@@
where $\mu_{0}$ is the mean of the initial population. It is relevant to notice that independently of the characteristic of $\Pi(f)$, the mean always halves at each generation. This result is simply understood considering that for each cell that inherits a fraction $f^{\prime}$ of the compound, its sibling inherits a fraction $1-f^{\prime}$. Hence, at every division, the compound count is diluted by an average factor of 2 in the population, and partition distributions with different properties lead to the same behavior of $\mu_{g}$ (see Figure 1(f)). Therefore, no relevant insights on the $\Pi(f)$ can be obtained from the first moment. We thus moved to consider the second moment, i.e., the variance. Similarly to what is done for the mean, the variance of the inherited fraction of elements can be expressed as:
@@
and in turn, the variance of a generation, starting from the consideration of the variance of a mixture of distributions, can be written as (see Supplementary materials for details):
@@
It is possible to obtain a recursive equation for $A_{g}$ (see Supplementary Notes), $A_{g}=A_{g-1}E[f^{2}]=A_{0}E[f^{2}]^{g}$ which leads to a compact form for Equation 8:
@@
Finally, with Equation 10, we can link the variance of the components distribution for the entire cell population at generation g, to the properties of $\Pi(f)$. It is interesting to highlight that there is no dependence on the bias asymmetry of the process. The distribution $\Pi(f)$ is necessarily symmetric, as daughter cells are indistinguishable. Therefore, by bias we refer to a situation where the most probable outcome of the division process is asymmetric. To this order, the only relevant feature shaping the dynamics is the extent of the fluctuations, regardless of their origin. Indeed, as shown in Figure 1(e), the same $\Pi(f)$, but with a different variance, shows a different behavior for $\sigma_{g}^{2}$.
Our analytical results show that the dynamical behavior of the population variance depends non-trivially on the second moment of the underlying partitioning distribution. The next section will be devoted to demonstrating how this condition can be exploited in experiments to measure the partition noise component and to evaluate the stability of our assumptions with respect to the concurrence of the other noise sources (see also Figure 1g).
@@ II.2 Measuring fluctuations via flow cytometry
As shown by Peruzzi et al. peruzzi2021asymmetric for leukemia cells growing in suspension, it is possible to use flow cytometry to follow the evolution of properly marked cellular elements in time. Here, we generalized the protocol considering a panel of colon cell lines, i.e. Caco2, a human epithelial colorectal adenocarcinoma cell line; HCT116, a human colorectal carcinoma cell line; and CCD-18Co, a normal human colon fibroblast cell line. To measure the degree of asymmetric division, we marked cells’ cytoplasm via a fluorescent dye and followed its dilution through successive generations. As explained in detail in the Materials and Methods section, the experimental procedure we used is based on the following main steps: staining of the selected compound, sorting of the initial population, plating, and acquisition of the distinct populations in the following days. The sorting allows us to obtain an initial population with a narrow peak at a specific fluorescence intensity, and it plays a dual role (a extended discussion can be found in the Supplementary Information): it indirectly induces cell synchronization and enables a clearer reconstruction of the coexisting generations.
To achieve this, we used fluorescence-activated cell sorting (FACS) to isolate a specific cell population based on cell morphology and fluorescence intensity. The sorted population is divided into distinct wells, one per acquisition, in equal numbers and kept under identical growth conditions to mitigate inoculum-dependent variability among samples  EnricoBena2021 . Figure 2(a) shows an example of the time course dynamics obtained with HCT116 cells (see Methods for details). Each distribution represents a different acquisition of a whole well at increasing times from plating (0 to 84 hours from the bottom to the top). It is possible to neatly observe the succession of peaks associated with different generations that succeed one another and the average decrease of the fluorescence intensity. Note that different generations can coexist at the same time point. One can spot the flow of cells towards higher generations and the behavior of the peak heights, which grow over time, reach dominance, and then decay. The shift on the y-axis allows us to follow the evolutionary dynamics while being able to see the overlapping of the same generation at different time points. As in each measurement, the distribution of fluorescent intensity, $I$, is a mixture of distributions corresponding to different generations:
@@
To obtain the necessary information on the properties of each generation, we performed a fitting of the data via a Gaussian Mixture Model (GMM) combined with an Expectation Maximization algorithm (details are explained in Materials and Methods). Indeed, assuming that each generation is log-normally distributed with mean $\mu_{g}$ and variance $\sigma_{g}^{2}$ of the corresponding Gaussian distribution, Equation 11 can be rewritten as :
@@
An example of the outcome of the GMM algorithm is shown in Figure 2(b) for two acquisitions (48h and 60 h from plating). Blue-shaded distributions represent experimental data, while the best solution of the Gaussian mixture is shown with a solid line. Each generation’s fit is identified with a different color. Upon an overnight, we can identify generations 0, 1, and 2 in both acquisitions. One can see that their relative fractions changed: generation 3 became dominant overnight, and generation 4 appeared.
@@
Figure 2(c) shows the behaviors of $\mu_{g}$ and $\sigma_{g}$ as functions of the generation number $g$ for three independent replicas of the time course experiments with CaCo2 cells. Dashed lines represent the best fit of the data against Equation 5 for the mean and Equation 10, which instead yields the variance $\sigma_{f}$ of the partition distribution probability, $\Pi(f)$. The outcomes of the fitting procedure on the time course dynamic for all the replicas of the three studied cell lines are reported in Figure 2(d). To evaluate the degree of asymmetry, we have defined the division asymmetry as the percentage of the coefficient of variation of the obtained partition distribution (CV: $\frac{\sigma}{\mu}$). In particular, Caco2 and CCD18Co cells show a higher level of asymmetry with respect to HCT116. As already stated in the description of the theoretical model, we consider only partitioning noise and neglect other potential sources: (i) production and degradation processes, and (ii) variability in cell cycle length. Since we use non-endogenous live fluorescent dyes, production processes are excluded by definition, while degradation results in a negligible time-dependent decrease in mean fluorescence intensity as the used dyes are optimized to maintain stable fluorescence levels for the duration of the experiment. Indeed, by looking at the data in Figure 2 (c), the values of the measured $\mu_{g}$ seem to only fluctuate around the theoretical line, but with the same slope, indicating that degradation can be neglected without major consequences. The effects of correlations between the inherited fractions of cellular components and the fluctuations in the duration of the cell cycle are, in principle, more difficult to disentangle. In this respect, it is widely known that cells adjust their growth/division strategy to mitigate size fluctuations at division miotto2024size ; thomas2018analysis . According to the ‘sizer’ model, they divide upon reaching a certain size so that smaller cells, i.e., those that inherit a smaller fraction of their mother’s volume, will have longer division times and vice versa. This results in fluctuations in the cell division times that may be coupled with partitioning noise.
@@
To assess the robustness of our model under such conditions, we performed simulations assuming that cells follow a sizer-like division strategy, while the marked component partition perfectly matches the cell size partition (see Supplementary Information for more details). This scenario maximizes the coupling between the two noise sources. As can be observed in Figure 1(g), the behaviors of the mean value $\mu_{g}$ and the variance $\sigma_{g}$ deviate from the case in which division time is uncoupled with partition noise. However, the values obtained are still within the estimated experimental error (see Supplementary Figure 3). It is worth noticing that the presence/absence of the increase in the mean, $\mu_{g}$, for newer generations may highlight the presence of the coupling between the two noise sources.